OpenStudy (anonymous):

solve the polynomial equation (x-3)^2=36

OpenStudy (curry):

x^2-6x= 27

OpenStudy (curry):

x^2 - 6x -27

OpenStudy (curry):

x+3 x-9 dont forget medal

OpenStudy (anonymous):

(x-3) = +sqrt36 or (x-3) = -sqrt36 x - 3 = 6 x = 9 x - 3 = -6 x = -3

OpenStudy (anonymous):

Solve x? FOIL (x - 3)^2 = (x - 3) (x - 3) = x^2 - 3x - 3x + 9 = 36 Square root the x^2 and the 36 That leaves you with x - 6x + 9 = 6 Combine like terms -5x + 9 = 6 Subtract 9 on both sides -5x = -3 Divide by -5 on both sides

OpenStudy (anonymous):

That leaves you with -3/-5 which = 3/5!!!

OpenStudy (anonymous):

Does that sound correct?

OpenStudy (radar):

Will the 3/5 check out? (3/5-3)^2=36 (-12/5)^2=36 144/25=36 Not looking that is a bad answer.

OpenStudy (anonymous):

So how do you do it then? :)

OpenStudy (radar):

Curry above worked it out for you. He may have been to brief but I will redo his work with explanations. First he expanded the expression enclosed in the parenthesis (x-3)^2 (x-3)(x-3)=x^2-6x+9=36 He then subtracted 9 from both sides getting: x^2-6x=27 He then subtracted 27 from both sides (leaving 0 on the right) x^2-6x-27=0 Now it is a quadratic in standard form. It is also factorable (x-9)(x+3)=0 Now it can be solved as the factors must equal 0 since the product is 0 x-9=0 ......... x must equal 9 x=9 is one root. x+3=0 ..........x must equal -3, x=-3 the second root. Do you follow this with understanding?